OpenStudy (anonymous):

For the functions below that have a removable discontinuity at x=a [if the function does not have a removable discontinuity, type in "n" below], state the value of g(a), where g(x) agrees with f(x) for x=a and is continuous everywhere.

OpenStudy (anonymous):

\[f(x)=\frac{ x^2-2x-8 }{ x+2} a=-2\]

OpenStudy (anonymous):

f(x)= \[\frac{ 3-\sqrt{3} }{ 9-x }\]

OpenStudy (anonymous):

a=9

OpenStudy (anonymous):

I got the first one

OpenStudy (anonymous):

the second one , no cule